xi (2,2) | (0.10) Consider the infinite repeated version of the game discount factor δ for which each player obtaining an average payoff around 5 is sustained by a SPNE strategy.
Answer #1

Looking at the game matrix and solving by the standard Nash equilibrium rules, we can see that Y strictly dominates X for both the players. Thus SPNE is (Y,Y) and payoff to each player is 1.

If instead we consider an infinitely repeated game with a discount factor delta and average it to 5,we get:

1+5 (18)2 (1+5)3

This is because discounted present value of any investment is given as -

PV= (1+r) where PV = present value, FV = future value, r = discount rate and n = no. of years.

Considering infinitely repeated game, the right side forms a GP.

Sum of infinite GP = rac{a}{1-r}

where a = 1 and r = 1+ 5 in our case.

Thus Sum = 1+δ

Thus 1+d

Rightarrow delta = rac{1}{4} = 25%